# Quantifying the Impact of the Si/O Interface in CCSN Explosions Using the Force Explosion Condition

### v1

@@ 4.2 Sensitivity to the density drop at the Si/O interface
Another way to observe the interplay between (parametrized) neutrino-driven convection and the accretion of the Si/O interface is to study how the explosion and the FEC+ change when only the Si/O interface changes, with everything else being the same. Therefore, we artificially smoothed the density drop at the Si/O interface of two different progenitors. The first two panels of Figures 6 and 7 show the smoothing of a $20M_{\odot}$ MESA progenitor (Farmer et al., 2016) and a $21M_{\odot}$ KEPLER progenitor (Sukhbold et al., 2016), respectively.
@@
The results of the simulations for the 20 $M_{\odot}$ MESA progenitor from Farmer et al. (2016) and for the 21 $M_{\odot}$ KEPLER progenitor from Sukhbold et al. (2016) are shown in Figures 6 and 7, respectively. These two progenitors were chosen because of their large interfaces and the number of computational zones within the interface, allowing for a more accurate smoothing procedure. Moreover, the accretion of the Si/O interface occurs at $\sim 0.22$ s after bounce for the 20 $M_{\odot}$ MESA progenitor, and at $0.1$ s post bounce for the 21 $M_{\odot}$ KEPLER progenitor. This allows us to study how the time of accretion changes the impact of the interface on the explosion.
The second panel of Figures 6 & 7 shows the average mass accretion rate in the gain region and, as expected, a smoother density profile at the Si/O interface corresponds to a shallower drop in mass accretion rate, which therefore prevents the shock from expanding. The third panels of those figures show the shock evolution for progenitors with different degrees of smoothing. The original progenitors explode and as the smoothing increases the shock expansion caused by the accretion of the Si/O interface becomes smaller, until eventually, it is not large enough to trigger explosion. This can be seen by looking at the last panel, which shows that for the unmodified progenitor, the FEC+ goes above the threshold as a consequence of the accretion of the Si/O interface, resulting in an explosion. As the smoothing increases, the increase is smaller and smaller until the FEC+ is not pushed above the threshold, and therefore the star does not explode.
The 20 $M_{\odot}$ MESA progenitor and the 21 $M_{\odot}$ KEPLER progenitor show the same general behavior concerning smoothing, and the threshold of the FEC+ is at around 0.28–0.3, consistent with the one found in the simulations discussed in sections 3 and 4.1. Moreover, the 21 $M_{\odot}$ KEPLER progenitor shows an early accretion of the Si/O interface, much like the 19.5 $M_{\odot}$ progenitor shown in Figure 2, the difference being that the Si/O interface for the 21 $M_{\odot}$ progenitor has a larger $\delta\rho/\rho$, and therefore explodes early. However, as the degree of smoothing increases, one can see that for $\Delta R/\Delta R_{\rm orig}\approx 10$ (i.e. the orange curve in Figure 7) the 21 $M_{\odot}$ progenitor has a very similar behavior to the 19.5 $M_{\odot}$ progenitor. An early accretion of a relatively large Si/O interface leads to the shock stalling at a large radius, and the FEC+ being pushed just below threshold. Eventually, the FEC+ slowly increases and the star explodes a few hundred milliseconds later. If one increases the smoothing further, the star does not explode.
@@ 5 Correlations between Si/O interface and FEC+
Similar conclusions can be drawn from the middle panel, in which $\Delta{\rm FEC+}$ (i.e. the difference between the FEC+ at $t=t_{\rm accr}^{\rm end}$ and at $t=t_{\rm accr}^{\rm start}$) is plotted against the value of the FEC+ at $t=t_{\rm accr}^{\rm start}$. The correlation between these two qunatities is quite evident. The reason is that, if the accretion occurs early in the post-bounce phase, the shock is still expanding, and therefore the FEC+ is also increasing. Therefore, one expects larger $\Delta{\rm FEC+}$ since not only does the accretion increases the FEC+, but the secular shock expansion also contributes to the increase. This is confirmed by the fact that simulations with early accretions (i.e. $t_{\rm accr}\lesssim 0.2$ s) also exhibit larger $\Delta{\rm FEC+}$. As mentioned above, the reason is that if the accretion of the Si/O interface occurs early enough in the post-bounce phase, then the FEC+ increases not only because of the accretion but also because of the secular shock expansion. Notice that this discussion is not valid for failed explosions and high-compactness progenitors.
@@
Finally, there are a few exceptions worth discussing. The crosses in Figure  8 are progenitors characterized by early accretion of the Si/O interface, i.e. $t_{\rm accr}<0.1$ s. Their shock radius and FEC+ evolution are very similar to the 19.5 $M_{\odot}$ progenitor shown in Figure 2. In other words, they are part of scenario 3) defined in Section 3, for which the early accretion of the Si/O interface does not push the FEC+ above the threshold, but instead pushes it very close to it, facilitating the occurrence of an explosion a few hundred milliseconds later. Therefore one expects the FEC+at $t=t_{\rm accr}^{\rm end}$ to be below the threshold, and the $\Delta{\rm FEC+}$ to be very large since the earlier the accretion, the faster the shock expansion, and therefore the larger the FEC+ increase. This is indeed confirmed by the left and middle panels in Figure 8. Most of the simulations shown as rainbow-colored dots instead belong to scenario 2) defined in Section 3. These are cases where the accretion of a large Si/O interface occurs during the stalled shock phase, pushes the FEC+ above the threshold, and therefore causes the explosion. As seen in the left panel of Figure 8, for some of them the FEC+ at $t=t_{\rm accr}^{\rm end}$ is however below the threshold, and $t_{\rm accr}>0.1$ s. These can be thought of as intermediate cases between scenarios 2) and 3). The accretion of the Si/O interface occurs at the very beginning of the stalled shock phase.
@@ 6 Conclusions
We also studied how the FEC+ (and therefore the explosion) varies depending on the strength of $\nu$-driven convection (i.e. the mixing length parameter $\alpha_{\rm MLT}$). We again showed that the explosion occurs only when the FEC+ goes above the threshold of 0.28–0.3 empirically derived in Section 3, which confirms the robustness of the FEC+ as an explosion condition. We showed that the value of the FEC+ at $t=t_{\rm accr}^{\rm start}$, as well as the change in FEC+ caused by the accretion of the Si/O interface, are positively correlated. They also positively correlate with $\alpha_{\rm MLT}$, the radius of the stalled shock and the increase in shock radius caused by the accretion of the interface. This confirms the key role that convection has in aiding the explosion, which is extremely important since it has been shown that, for example, three-dimensional asymmetries in the pre-SN progenitor (Müller et al., 2016b; Yadav et al., 2020; Couch et al., 2015; Fields & Couch, 2021) can significantly change the strength of convection that develops in the post-bounce phase. Moreover, it is still unclear if and how numerical resolution (Radice et al., 2016; Abdikamalov et al., 2015) or the dimensionality of the problem (Couch & Ott, 2015; Murphy et al., 2013) can artificially alter the efficiency of convection.
@@
Finally, we analyzed how the change in FEC+ is related to the value of the FEC+ right before accretion of the Si/O interface through the shock, and also to the density drop at the Si/O interface $\delta\rho_{\rm Si/O}^{2}/\rho_{\rm Si/O}^{2}$. We concluded that for the accretion of the Si/O interface to lead to successful explosions, the FEC+ should already be close to the explosion threshold, depending on how large the density drop at the Si/O interface is. We then quantified the effect of the Si/O interface accretion by analyzing the exploding progenitors with $\xi_{2.0}\lesssim 0.5$.
The analysis for progenitors accreting the interface early (i.e. before $\sim 0.18$ s after bounce) is complicated by the fact that the FEC+ changes not only because of the shock expansion caused by the accretion of the Si/O interface, but also because of a secular shock expansion. By focusing only on progenitors where the secular shock expansion is negligible, we find that the accretion of the Si/O interface can contribute between 5 % and 15% to the overall explosion condition, i.e. between $20\%$ and $50\%$ of the overall effect of convection, which has been estimated to decrease the explosion condition by about 2530 % (Gogilashvili et al., 2023a).

### v2

@@ 4.2 Sensitivity to the density drop at the Si/O interface
Another way to observe the interplay between (parametrized) neutrino-driven convection and the accretion of the Si/O interface is to study how the explosion and the FEC+ change when only the Si/O interface changes, with everything else being the same. Therefore, we artificially smoothed the density drop at the Si/O interface of two different progenitors. The first two panels of Figures 9 and 10 show the smoothing of a $20M_{\odot}$ MESA progenitor (Farmer et al., 2016) and a $21M_{\odot}$ KEPLER progenitor (Sukhbold et al., 2016), respectively.
@@
The results of the simulations for the 20 $M_{\odot}$ MESA progenitor from Farmer et al. (2016) and for the 21 $M_{\odot}$ KEPLER progenitor from Sukhbold et al. (2016) are shown in Figures 9 and 10, respectively. These two progenitors were chosen because of their large interfaces and the number of computational zones within the interface, allowing for a more accurate smoothing procedure. Moreover, the accretion of the Si/O interface occurs at $\sim 0.22$ s after bounce for the 20 $M_{\odot}$ MESA progenitor, and at $0.1$ s post bounce for the 21 $M_{\odot}$ KEPLER progenitor. This allows us to study how the time of accretion changes the impact of the interface on the explosion.
The second panel of Figures 9 & 10 shows the average mass accretion rate in the gain region and, as expected, a smoother density profile at the Si/O interface corresponds to a shallower drop in mass accretion rate, which therefore prevents the shock from expanding. The third panels of those figures show the shock evolution for progenitors with different degrees of smoothing. The original progenitors explode and as the smoothing increases the shock expansion caused by the accretion of the Si/O interface becomes smaller, until eventually, it is not large enough to trigger an explosion. This can be seen by looking at the last panel, which shows that for the unmodified progenitor, the FEC+ goes above the threshold as a consequence of the accretion of the Si/O interface, resulting in an explosion. As the smoothing increases, the increase is smaller and smaller until the FEC+ is not pushed above the threshold, and therefore the star does not explode.
The 20 $M_{\odot}$ MESA progenitor and the 21 $M_{\odot}$ KEPLER progenitor show the same general behavior concerning smoothing, and the threshold of the FEC+ is at around 0.28–0.3, consistent with the one found in the simulations discussed in sections 3 and 4.1. Moreover, the 21 $M_{\odot}$ KEPLER progenitor shows an early accretion of the Si/O interface, much like the 19.5 $M_{\odot}$ progenitor shown in Figure 2, the difference being that the Si/O interface for the 21 $M_{\odot}$ progenitor has a larger $\delta\rho/\rho$, and therefore explodes early. However, as the degree of smoothing increases, one can see that for $\Delta R/\Delta R_{\rm orig}\approx 10$ (i.e. the orange curve in Figure 10) the 21 $M_{\odot}$ progenitor has very similar behavior to the 19.5 $M_{\odot}$ progenitor. An early accretion of a relatively large Si/O interface leads to the shock stalling at a large radius, and the FEC+ being pushed just below threshold. Eventually, the FEC+ slowly increases and the star explodes a few hundred milliseconds later. If one increases the smoothing further, the star does not explode.
@@ 5 Correlations between Si/O interface and FEC+
Similar conclusions can be drawn from the middle panel, in which $\Delta{\rm FEC+}$ (i.e. the difference between the FEC+ at $t=t_{\rm accr}^{\rm end}$ and at $t=t_{\rm accr}^{\rm start}$) is plotted against the value of the FEC+ at $t=t_{\rm accr}^{\rm start}$. The correlation between these two quantities is quite evident. The reason is that, if the accretion occurs early in the post-bounce phase, the shock is still expanding, and therefore the FEC+ is also increasing. Therefore, one expects larger $\Delta{\rm FEC+}$ since not only does the accretion increase the FEC+, but the secular shock expansion also contributes to the increase. This is confirmed by the fact that simulations with early accretions (i.e. $t_{\rm accr}\lesssim 0.2$ s) also exhibit larger $\Delta{\rm FEC+}$. As mentioned above, the reason is that if the accretion of the Si/O interface occurs early enough in the post-bounce phase, then the FEC+ increases not only because of the accretion but also because of the secular shock expansion. Notice that this discussion is not valid for failed explosions and high-compactness progenitors.
@@
Finally, there are a few exceptions worth discussing. The crosses in Figure  11 are progenitors characterized by early accretion of the Si/O interface, i.e. $t_{\rm accr}<0.1$ s. Their shock radius and FEC+ evolution are very similar to the 19.5 $M_{\odot}$ progenitor shown in Figure 2. In other words, they are part of scenario 3) defined in Section 3, for which the early accretion of the Si/O interface does not push the FEC+ above the threshold, but instead pushes it very close to it, facilitating the occurrence of an explosion a few hundred milliseconds later. Therefore one expects the FEC+at $t=t_{\rm accr}^{\rm end}$ to be below the threshold, and the $\Delta{\rm FEC+}$ to be very large since the earlier the accretion, the faster the shock expansion, and therefore the larger the FEC+ increase. This is indeed confirmed by the left and middle panels in Figure 11. Most of the simulations shown as rainbow-colored dots instead belong to scenario 2) defined in Section 3. These are cases where the accretion of a large Si/O interface occurs during the stalled shock phase, pushes the FEC+ above the threshold, and therefore causes the explosion. As seen in the left panel of Figure 11, for some of them the FEC+ at $t=t_{\rm accr}^{\rm end}$ is however below the threshold, and $t_{\rm accr}>0.1$ s. These can be thought of as intermediate cases between scenarios 2) and 3). The accretion of the Si/O interface occurs at the very beginning of the stalled shock phase.
@@ 6 Conclusions
We also studied how the FEC+ (and therefore the explosion) varies depending on the strength of $\nu$-driven convection (i.e. the mixing length parameter $\alpha_{\rm MLT}$). We again showed that the explosion occurs only when the FEC+ goes above the threshold of 0.28–0.3 empirically derived in Section 3, which confirms the robustness of the FEC+ as an explosion condition. We showed that the value of the FEC+ at $t=t_{\rm accr}^{\rm start}$, as well as the change in FEC+ caused by the accretion of the Si/O interface, are positively correlated. They also positively correlate with $\alpha_{\rm MLT}$, the radius of the stalled shock, and the increase in shock radius caused by the accretion of the interface. This confirms the key role that convection has in aiding the explosion, which is extremely important since it has been shown that, for example, three-dimensional asymmetries in the pre-SN progenitor (Müller et al., 2016b; Yoshida et al., 2019; Yadav et al., 2020; Couch et al., 2015; Fields & Couch, 2021) can significantly change the strength of convection that develops in the post-bounce phase. Moreover, it is still unclear if and how numerical resolution (Radice et al., 2016; Abdikamalov et al., 2015) or the dimensionality of the problem (Couch & Ott, 2015; Murphy et al., 2013) can artificially alter the efficiency of convection.
@@
Finally, we analyzed how the change in FEC+ is related to the value of the FEC+ right before the accretion of the Si/O interface through the shock, and also to the density drop at the Si/O interface $\delta\rho_{\rm Si/O}^{2}/\rho_{\rm Si/O}^{2}$. We concluded that for the accretion of the Si/O interface to lead to successful explosions, the FEC+ should already be close to the explosion threshold, depending on how large the density drop at the Si/O interface is. We then quantified the effect of the Si/O interface accretion by analyzing the exploding progenitors with $\xi_{2.0}\lesssim 0.5$.
The analysis for progenitors accreting the interface early (i.e. before $\sim 0.18$ s after bounce) is complicated by the fact that the FEC+ changes not only because of the shock expansion caused by the accretion of the Si/O interface but also because of a secular shock expansion. By focusing only on progenitors where the secular shock expansion is negligible, we find that the accretion of the Si/O interface can contribute between 5 % and 15% to the overall explosion condition, i.e. between $20\%$ and $50\%$ of the overall effect of convection, which has been estimated to decrease the explosion condition by about 2530 % (Gogilashvili et al., 2023a).